# Corona rigidity

### v3

@@ Abstract.
We give a unified overview of the study of the effects of additional set theoretic axioms on quotient structures. Our focus is on rigidity, measured in terms of existence (or rather non-existence) of suitably non-trivial automorphisms of the quotients in question. A textbook example for the study of this topic is the Boolean algebra $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$, whose behavior is the template around which this survey revolves: Forcing axioms imply that all of its automorphisms are trivial, in the sense that they are induced by almost permutations of ${\mathbb{N}}$, while under the Continuum Hypothesis this rigidity fails and $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ admits uncountably many non-trivial automorphisms. We consider far-reaching generalizations of this phenomenon and present a wide variety of situations where analogous patterns persist, focusing mainly (but not exclusively) on the categories of Boolean algebras, Čech–Stone remainders, and $\mathrm{C}^{*}$-algebras. We survey the state of the art and the future prospects of this field, discussing the major open problems and outlining the main ideas of the proofs whenever possible.
The third author was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure, through SFB 1442, by the ERC Advanced Grant 834267 - AMAREC, and by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 891709. The fourth author was supported by an Emergence en recherche Idex grant from Université Paris Cité.
@@ 1. Introduction
Fast-forwarding 30 years, the same pattern was isolated again in a different quotient structure, this time originating in the context of algebras of operators on Hilbert spaces (see §4.1 for the definitions): The Calkin algebra. The Calkin algebra $\mathcal{Q}(H)$ is the quotient of the $\mathrm{C}^{*}$-algebra of all bounded linear operators on a complex, infinite-dimensional, separable Hilbert space $H$, by the ideal of all compact operators. In this case, we say that an automorphism of $\mathcal{Q}(H)$ is inner if it is implemented by a unitary in $\mathcal{Q}(H)$ (equivalently, it is lifted by a conjugation by an isometry between two closed subspaces of $H$ with finite co-dimension). In the early 20th century Weyl and von Neumann initiated the study of unitary equivalence of self-adjoint operators modulo compact perturbations. Their elegant characterization of this equivalence was extended to normal operators ([10, 152]) and then to essentially normal operators, those whose images in $\mathcal{Q}(H)$ are normal ([20]). While two normal operators are unitarily equivalent modulo compact perturbations if and only if their images in the Calkin algebra are conjugate by an automorphism, it was not clear whether this conclusion could be extended to essentially normal operators. In 1973 (a few years before Shelah’s result) Brown, Douglas, and Fillmore asked whether $\mathcal{Q}(H)$ has an outer (i.e. not inner) automorphism. If one replaces ‘trivial automorphism’ with ‘inner automorphism’ then the exact analogs of Rudin’s and Shelah’s results hold for $\mathcal{Q}(H)$: Phillips and Weaver ([124]) showed that CH implies that $\mathcal{Q}(H)$ admits $2^{\mathfrak{c}}$ automorphisms (most of them outer), while the first author ([58]) proved that under forcing axioms all of its automorphisms are inner.This implies that, consistently with ZFC, two essentially normal operators are unitarily equivalent modulo compact perturbations if and only if their images in the Calkin algebra are conjugate by an automorphism. It remains open whether this is a theorem of ZFC.
@@
In §2 we present the abstract rigidity problem. §3 serves as a warm-up. Here we present some basic examples of quotient structures from the worlds of Boolean algebras and of topological spaces, in the form of Čech–Stone remainders. In §4 we introduce $\mathrm{C}^{*}$-algebras as well as the framework which eventually motivated and ignited the study of rigidity questions in the noncommutative setting. §5 is devoted to Ulam-stability and to the problem of obtaining algebraically trivial liftings from topologically trivial ones. [^3] An isomorphism is called topologically trivial if it has a Borel lifting, see Definition 2.2. This topic, as interesting and as necessary for understanding quotient rigidity, has little to do with logic and readers interested only in logical aspects may want to skip it. §6 and §7 are the main body of this survey, and they portray two faces of the same coin. The former describes various constructions of isomorphisms of Borel quotient structures which are not topologically trivial, under set-theoretic assumptions such as CH and some of its weakenings. The latter, on the other hand, presents rigidity results, either obtained via forcing or as a consequence of Forcing Axioms, and instances where all automorphisms of certain quotient structures are topologically trivial. In §8 we explore rigidity results for maps more general than isomorphisms and automorphisms, such as endomorphisms of the Calkin algebra. §9 is devoted to large quotients of the form $\mathcal{P}(\kappa)/\operatorname{Fin}$ and $\mathcal{Q}(\ell_{2}(\kappa))$ for uncountable $\kappa$, while §10 focuses on rigidity results on coronas of uniform Roe algebras, $\mathrm{C}^{*}$-algebras arising from coarse metric spaces. In §11 we provide a quick overview of a selection of topics loosely related to the contents of this survey. Finally §12 is reserved for remarks of metamathematical nature on the rigidity problems discussed in the rest of the survey.
@@ 2. The general rigidity question
In this section we introduce a unifying framework for describing all the categories and rigidity instances considered in the rest of the survey. Readers interested in less abstract results are encouraged to skip ahead and refer to this section as needed. Two critically important notions are that of topologically trivial and algebraically trivial automorphisms (respectively Definition 2.2 and 2.3), before which we isolate the class of quotients we aim to study, with a definition that generalizes both the Boolean algebra $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ and the Calkin algebra $\mathcal{Q}(H)$.
@@
    The Continuum Hypothesis (CH) implies that ${\mathcal{M}}/E$ has $2^{\mathfrak{c}}$ automorphisms (and therefore $2^{\mathfrak{c}}$ topologically nontrivial automorphisms). It also implies that ${\mathcal{M}}/E$ and ${\mathcal{N}}/F$ are isomorphic, unless there is an obvious obstruction for this. [^6] A reader familiar with Woodin’s $\Sigma^{2}_{1}$ absoluteness theorem will recognize its relevance; more on this in §12.
@@
A simpler proof of Truss’s theorem used in the previous example was announced in [1]. There is unfortunately a gap in this proof; see [36, Remark 11.A.6].
@@ 3. The abelian case: Boolean algebras and Čech–Stone remainders
We now turn to two classes of Borel quotient structures that generalize $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ in different directions:
@@
Rigidity problems for such quotients can be traced back at least to [47, p. 38-39], where Erdös and Ulam asked whether the quotients over the two ideals, known as the asymptotic density zero ideal ${\mathcal{Z}}_{0}$ and the logarithmic density zero ideal ${\mathcal{Z}}_{\log}$, are isomorphic, where
@@
The unexpected (at the time) answer to this question will be discussed in §6.1.2 and §7.2, where we will have something interesting to say also for the classes of analytic P-ideals and of generalized density ideals, to which both ${\mathcal{Z}}_{0}$ and ${\mathcal{Z}}_{\log}$ belong, and which we briefly anticipate here.
@@
Let ${\mathbb{N}}=\bigsqcup_{m}I_{m}$ be a partition of ${\mathbb{N}}$ into finite sets and $\mu_{m}$ be a submeasure concentrated on $I_{m}$ for all $m$. Consider the measure $\mu(X)=\sup_{m}\mu_{m}(X\cap I_{m})$. An ideal is a generalized density ideal if it is of the form $\operatorname{Exh}(\mu)$ for a $\mu$ arising this way. If each $\mu_{n}$ is a measure, then this is a density ideal.
Density ideals were introduced in [50, Definition 1.13.1], and it is not difficult to see that both ${\mathcal{Z}}_{0}$ and ${\mathcal{Z}}_{\log}$ are density ideals ([50, Theorem 1.13.3 (a)]; for a fix to part (b) of this theorem see [162]). We also have the following results of Mazur ([108]) and Solecki ([154]).
@@
(The last line is the easy [50, Lemma 1.2.3].) Note that if ${\mathcal{I}}$ is not an $F_{\sigma}$ ideal, then there is a decreasing sequence of ${\mathcal{I}}$-positive sets $A_{n}$ such that $\lim_{n}\mu(A_{n})=0$. This shows a failure of the countable saturation of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ as a discrete structure, and we shall see why this is relevant to us in §6.1.2.
@@
Given a locally compact space $X$, its Čech–Stone compactification $\beta X$ is the ‘largest’ compact space in which $X$ sits densely as an open set. $\beta X$ is defined as the space with the universal property that if $K$ is a compact Hausdorff space and $f\colon X\to K$ is continuous, one can find a continuous extension $\beta f\colon\beta X\to K$ [^9] Functional analyst’s definition of $\beta X$: It is the Gelfand spectrum of the $\mathrm{C}^{*}$-algebra $C_{b}(X)$ of all continuous, bounded, complex-valued functions on $X$ (see §4.1 and §4.2). The compact space $\beta X\setminus X$ is the Čech–Stone remainder (or the corona) of $X$. Čech–Stone remainders of zero-dimensional spaces and their homeomorphisms from a set theoretic point of view were analyzed in [50] with the aid of Boolean algebraic methods.
@@ 4.1. Preliminaries
Alternatively, $\mathrm{C}^{*}$-algebras can be defined by providing an axiomatization in the following way. A -Banach algebra is a tuple $(A,\|\cdot\|,+,\cdot,^{*})$ such that $(A,+,\cdot)$ is a $\mathbb{C}$-algebra, the tuple $(A,\|\cdot\|,+)$ is a complex Banach space, is a linear involution, and the following conditions hold:
@@
Albeit seemingly mild, the $\mathrm{C}^{*}$-equality is a powerful condition that tightly connects the topological nature of $\mathrm{C}^{*}$-algebras with their algebraic properties. It implies, for instance, that in a $\mathrm{C}^{*}$-algebra the norm is completely determined by its algebraic structure, and therefore it is unique ([59, Corollary 1.3.3]). It moreover entails that -homomorphisms between $\mathrm{C}^{*}$-algebras, namely linear multiplicative maps preserving the involution (the relevant arrows in this category), are automatically continuous ([59, Lemma 1.2.10]).
By the Gelfand–Naimark–Segal (GNS) construction (see [59, §1.10]) every $\mathrm{C}^{*}$-algebra $A$ is isometrically isomorphic to a concrete $\mathrm{C}^{*}$-algebra in some $\mathcal{B}(H)$. Equivalently, the GNS-construction shows that concrete $\mathrm{C}^{*}$-algebras are axiomatized by the conditions listed above, allowing us to identify the two definitions.
@@ 4.1.1. Examples I: Abelian $\mathrm{C}^{*}$-algebras
One of the cornerstones in the theory of $\mathrm{C}^{*}$-algebras is the Gelfand Representation Theorem ([59, Theorem 1.3.1]), which states that every abelian $\mathrm{C}^{*}$-algebra is isomorphic to one of the form $C_{0}(X)$, for some locally compact $X$, where
@@
is considered with the supremum norm and pointwise operations. The Gelfand Representation Theorem establishes a natural equivalence between the category of unital abelian $\mathrm{C}^{*}$-algebras and the category of compact Hausdorff spaces (e.g. [59, Theorem 1.3.2]). This duality, usually referred to as the Gelfand–Naimark Duality, motivates the leading philosophy in the subject, namely that $\mathrm{C}^{*}$-algebras provide the noncommutative analogue of topological spaces. Indeed, many definitions, techniques, and results in $\mathrm{C}^{*}$-algebras arise as adaptations of ideas that stem from topology.
@@ 4.1.2. Examples II: General $\mathrm{C}^{*}$-algebras
    Given a separable, infinite-dimensional Hilbert space $H$, the Calkin algebra is the quotient $\mathcal{Q}(H):=\mathcal{B}(H)/\mathcal{K}(H)$. Being the quotient of a $\mathrm{C}^{*}$-algebra by a two-sided norm-closed ideal, it is itself a $\mathrm{C}^{*}$-algebra (this is nontrivial—see [59, Lemma 2.5.2]) with the operations and norm induced by $\mathcal{B}(H)$.
@@
    if $A$ is unital, $\sigma(a):=\{\lambda\in\mathbb{C}\mid a-\lambda 1\text{ is invertible}\}$, [^10] The spectrum of a normal element $a$ of a non-unital $\mathrm{C}^{*}$-algebra $A$ is defined as the spectrum of $a$ in the unitization of $A$. is the spectrum of $a$,

### v4

@@ Abstract.
We give a unified overview of the study of the effects of additional set theoretic axioms on quotient structures. Our focus is on rigidity, measured in terms of existence (or rather non-existence) of suitably non-trivial automorphisms of the quotients in question. A textbook example for the study of this topic is the Boolean algebra $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$, whose behavior is the template around which this survey revolves: Forcing axioms imply that all of its automorphisms are trivial, in the sense that they are induced by almost permutations of ${\mathbb{N}}$, while under the Continuum Hypothesis this rigidity fails and $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ admits uncountably many non-trivial automorphisms. We consider far-reaching generalisations of this phenomenon and present a wide variety of situations where analogous patterns persist, focusing mainly (but not exclusively) on the categories of Boolean algebras, Čech–Stone remainders, and $\mathrm{C}^{*}$-algebras. We survey the state of the art and the future prospects of this field, discussing the major open problems and outlining the main ideas of the proofs whenever possible.
@@ 1. Introduction
Fast-forwarding 30 years, the same pattern was isolated again in a different quotient structure, this time originating in the context of algebras of operators on Hilbert spaces (see §4.1 for the definitions): The Calkin algebra. The Calkin algebra $\mathcal{Q}(H)$ is the quotient of the $\mathrm{C}^{*}$-algebra of all bounded linear operators on a complex, infinite-dimensional, separable Hilbert space $H$, by the ideal of all compact operators. In this case, we say that an automorphism of $\mathcal{Q}(H)$ is inner if it is implemented by a unitary in $\mathcal{Q}(H)$ (equivalently, it is lifted by a conjugation by an isometry between two closed subspaces of $H$ with finite co-dimension). In the early 20th century Weyl and von Neumann initiated the study of unitary equivalence of self-adjoint operators modulo compact perturbations. Their elegant characterisation of this equivalence was extended to normal operators ([10, 159]) and then to essentially normal operators, those whose images in $\mathcal{Q}(H)$ are normal ([20]). While two normal operators are unitarily equivalent modulo compact perturbations if and only if their images in the Calkin algebra are conjugate by an automorphism, it was not clear whether this conclusion could be extended to essentially normal operators. In 1973 (a few years before Shelah’s result) Brown, Douglas, and Fillmore asked whether $\mathcal{Q}(H)$ has an outer (i.e. not inner) automorphism. If one replaces ‘trivial automorphism’ with ‘inner automorphism’ then the exact analogs of Rudin’s and Shelah’s results hold for $\mathcal{Q}(H)$: Phillips and Weaver ([130]) showed that CH implies that $\mathcal{Q}(H)$ admits $2^{\mathfrak{c}}$ automorphisms (most of them outer), while the first author ([60]) proved that under forcing axioms all of its automorphisms are inner.This implies that, consistently with ZFC, two essentially normal operators are unitarily equivalent modulo compact perturbations if and only if their images in the Calkin algebra are conjugate by an automorphism. It remains open whether this is a theorem of ZFC.
@@
In §2 we present the abstract rigidity problem. §3 serves as a warm-up. Here we present some basic examples of quotient structures from the worlds of Boolean algebras and of topological spaces, in the form of Čech–Stone remainders. In §4 we introduce $\mathrm{C}^{*}$-algebras as well as the framework which eventually motivated and ignited the study of rigidity questions in the noncommutative setting. §5 is devoted to Ulam-stability and to the problem of obtaining algebraically trivial liftings from topologically trivial ones. [^3] An isomorphism is called topologically trivial if it has a Borel lifting, see Definition 2.2. This topic, as interesting and as necessary for understanding quotient rigidity, has little to do with logic and readers interested only in logical aspects may want to skip it. §6 and §7 are the main body of this survey, and they portray two faces of the same coin. The former describes various constructions of isomorphisms of Borel quotient structures which are not topologically trivial, under set-theoretic assumptions such as CH and some of its weakenings. The latter, on the other hand, presents rigidity results, either obtained via forcing or as a consequence of Forcing Axioms, and instances where all automorphisms of certain quotient structures are topologically trivial. In §8 we explore rigidity results for maps more general than isomorphisms and automorphisms, such as endomorphisms of the Calkin algebra. §9 is devoted to large quotients of the form $\mathcal{P}(\kappa)/\operatorname{Fin}$ and $\mathcal{Q}(\ell_{2}(\kappa))$ for uncountable $\kappa$, while §10 focuses on rigidity results on uniform Roe and Higson coronas, $\mathrm{C}^{*}$-algebras arising from coarse metric spaces. In §11 we provide a quick overview of a selection of topics loosely related to the contents of this survey. Finally §12 is reserved for remarks of metamathematical nature on the rigidity problems discussed in the rest of the survey.
@@ Acknowledgments
Funding
IF was supported by NSERC. AVa was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 –390685587, Mathematics Münster: Dynamics–Geometry–Structure, through SFB 1442, by the ERC Advanced Grant 834267 - AMAREC, and by the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 891709. AVi was supported by the Institut Universitaire de France.
@@ 2. The general rigidity question
In this section we introduce a unifying framework for describing all the categories and rigidity instances considered in the rest of the survey. Readers interested in less abstract results are encouraged to skip ahead and refer to this section as needed. Two critically important notions are that of topologically trivial and algebraically trivial automorphisms (respectively Definition 2.2 and 2.3), before which we isolate the class of quotients we aim to study, with a definition that generalises both the Boolean algebra $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ and the Calkin algebra $\mathcal{Q}(H)$.
@@
    The Continuum Hypothesis (CH) implies that ${\mathcal{M}}/E$ has $2^{\mathfrak{c}}$ automorphisms (and therefore $2^{\mathfrak{c}}$ topologically nontrivial automorphisms). It also implies that ${\mathcal{M}}/E$ and ${\mathcal{N}}/F$ are isomorphic, unless there is an obvious obstruction for this. [^6] A reader familiar with Woodin’s $\Sigma^{2}_{1}$ absoluteness theorem will recognise its relevance; more on this in §12.
@@
A simpler proof of Truss’s theorem used in the previous example was announced in [1]. There is unfortunately a gap in this proof; see [37, Remark 11.A.6].
@@ 3. The abelian case: Boolean algebras and Čech–Stone remainders
We now turn to two classes of Borel quotient structures that generalise $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ in different directions:
@@
Rigidity problems for such quotients can be traced back at least to [49, p. 38-39], where Erdös and Ulam asked whether the quotients over the two ideals, known as the asymptotic density zero ideal ${\mathcal{Z}}_{0}$ and the logarithmic density zero ideal ${\mathcal{Z}}_{\log}$, are isomorphic, where
@@
The unexpected (at the time) answer to this question will be discussed in §6.1.2 and §7.2, where we will have something interesting to say also for the classes of analytic P-ideals and of generalised density ideals, to which both ${\mathcal{Z}}_{0}$ and ${\mathcal{Z}}_{\log}$ belong, and which we briefly anticipate here.
@@
Let ${\mathbb{N}}=\bigsqcup_{m}I_{m}$ be a partition of ${\mathbb{N}}$ into finite sets and $\mu_{m}$ be a submeasure concentrated on $I_{m}$ for all $m$. Consider the measure $\mu(X)=\sup_{m}\mu_{m}(X\cap I_{m})$. An ideal is a generalised density ideal if it is of the form $\operatorname{Exh}(\mu)$ for a $\mu$ arising this way. If each $\mu_{n}$ is a measure, then this is a density ideal.
Density ideals were introduced in [52, Definition 1.13.1], and it is not difficult to see that both ${\mathcal{Z}}_{0}$ and ${\mathcal{Z}}_{\log}$ are density ideals ([52, Theorem 1.13.3 (a)]; for a fix to part (b) of this theorem see [169]). We also have the following results of Mazur ([114]) and Solecki ([161]).
@@
(The last line is the easy [52, Lemma 1.2.3].) Note that if ${\mathcal{I}}$ is not an $F_{\sigma}$ ideal, then there is a decreasing sequence of ${\mathcal{I}}$-positive sets $A_{n}$ such that $\lim_{n}\mu(A_{n})=0$. This shows a failure of the countable saturation of $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ as a discrete structure, and we shall see why this is relevant to us in §6.1.2.
@@
Given a locally compact space $X$, its Čech–Stone compactification $\beta X$ is the ‘largest’ compact space in which $X$ sits densely as an open set. $\beta X$ is defined as the space with the universal property that if $K$ is a compact Hausdorff space and $f\colon X\to K$ is continuous, one can find a continuous extension $\beta f\colon\beta X\to K$ [^9] Functional analyst’s definition of $\beta X$: It is the Gelfand spectrum of the $\mathrm{C}^{*}$-algebra $C_{b}(X)$ of all continuous, bounded, complex-valued functions on $X$ (see §4.1 and §4.2). The compact space $\beta X\setminus X$ is the Čech–Stone remainder (or the corona) of $X$. Čech–Stone remainders of zero-dimensional spaces and their homeomorphisms from a set theoretic point of view were analysed in [52] with the aid of Boolean algebraic methods.
@@ 4.1. Preliminaries
Alternatively, $\mathrm{C}^{*}$-algebras can be defined by providing an axiomatisation in the following way. A -Banach algebra is a tuple $(A,\|\cdot\|,+,\cdot,^{*})$ such that $(A,+,\cdot)$ is a $\mathbb{C}$-algebra, the tuple $(A,\|\cdot\|,+)$ is a complex Banach space, is a linear involution, and the following conditions hold:
@@
Albeit seemingly mild, the $\mathrm{C}^{*}$-equality is a powerful condition that tightly connects the topological nature of $\mathrm{C}^{*}$-algebras with their algebraic properties. It implies, for instance, that in a $\mathrm{C}^{*}$-algebra the norm is completely determined by its algebraic structure, and therefore it is unique ([61, Corollary 1.3.3]). It moreover entails that -homomorphisms between $\mathrm{C}^{*}$-algebras, namely linear multiplicative maps preserving the involution (the relevant arrows in this category), are automatically continuous ([61, Lemma 1.2.10]).
By the Gelfand–Naimark–Segal (GNS) construction (see [61, §1.10]) every $\mathrm{C}^{*}$-algebra $A$ is isometrically isomorphic to a concrete $\mathrm{C}^{*}$-algebra in some $\mathcal{B}(H)$. Equivalently, the GNS-construction shows that concrete $\mathrm{C}^{*}$-algebras are axiomatised by the conditions listed above, allowing us to identify the two definitions.
@@ 4.1.1. Examples I: Abelian $\mathrm{C}^{*}$-algebras
One of the cornerstones in the theory of $\mathrm{C}^{*}$-algebras is the Gelfand Representation Theorem ([61, Theorem 1.3.1]), which states that every abelian $\mathrm{C}^{*}$-algebra is isomorphic to one of the form $C_{0}(X)$, for some locally compact $X$, where
@@
is considered with the supremum norm and pointwise operations. The Gelfand Representation Theorem establishes a natural equivalence between the category of unital abelian $\mathrm{C}^{*}$-algebras and the category of compact Hausdorff spaces (e.g. [61, Theorem 1.3.2]). This duality, usually referred to as the Gelfand–Naimark Duality, motivates the leading philosophy in the subject, namely that $\mathrm{C}^{*}$-algebras provide the noncommutative analogue of topological spaces. Indeed, many definitions, techniques, and results in $\mathrm{C}^{*}$-algebras arise as adaptations of ideas that stem from topology.
@@ 4.1.2. Examples II: General $\mathrm{C}^{*}$-algebras
    Given a separable, infinite-dimensional Hilbert space $H$, the Calkin algebra is the quotient $\mathcal{Q}(H):=\mathcal{B}(H)/\mathcal{K}(H)$. Being the quotient of a $\mathrm{C}^{*}$-algebra by a two-sided norm-closed ideal, it is itself a $\mathrm{C}^{*}$-algebra (this is nontrivial—see [61, Lemma 2.5.2]) with the operations and norm induced by $\mathcal{B}(H)$.
@@
    if $A$ is unital, $\sigma(a):=\{\lambda\in\mathbb{C}\mid a-\lambda 1\text{ is invertible}\}$, [^10] The spectrum of a normal element $a$ of a non-unital $\mathrm{C}^{*}$-algebra $A$ is defined as the spectrum of $a$ in the unitisation of $A$. is the spectrum of $a$,